if f(x)=x^2, find f(5-sqrt2)
show work please

Answer by fractalier(6550)

If f(x) = x^2, then
f(5 - sqrt(2)) = (5 - sqrt(2))^2 =
(5 - sqrt(2)) * (5 - sqrt(2)) =
25 - 5*sqrt(2) - 5*sqrt(2) + 2 =
27 - 10*sqrt(2)